mikewwe13:

Use a system of equations to solve this problem. Bronze is a mix, or alloy, of tin and copper. A metal worker needs 100 g of bronze that is 25% tin. He has one tin/copper alloy that is 5% tin and another tin/copper alloy that is 45% tin. Let x = the number of grams of the 5% tin alloy. Let y = the number of grams of the 45% tin alloy. How many grams of each alloy should the metal worker combine? Enter your answers in the boxes. ______ g of the 5% tin alloy _____ g of the 45% tin alloy

mikewwe13:

@Vocaloid

Vocaloid:

total mass of metal = 100 x + y = 100 according to the variables defined in the problem 25% tin of 100g sample means 25g of tin tin from x + tin from y = 25g (0.05)x + (0.45)y = 25 solve the following system: x + y = 100 (0.05)x + (0.45)y = 25

mikewwe13:

25 and 50